OpenStudy (anonymous):

How would I figure out the equation of a line with the points (4,1) and (-3,-1)

OpenStudy (science0229):

First, you would want to find the slope of that line. For 2 points (a,b) and (c,d), the slope is \[\frac{ d-b }{ c-a }\]

OpenStudy (science0229):

So (-1-1)/(-3-4)=-2/-7=2/7 The slope for the line is 2/7.

OpenStudy (science0229):

Using the point-slope formula and the point (4,1), y-1=(2/7)(x-4) or y=(2/7)x-(1/7)

OpenStudy (anonymous):

where did the 1/7 come from

OpenStudy (anonymous):

@science0229

OpenStudy (jdoe0001):

hmmm do you know how to get the slope of the points -> (4,1) and (-3,-1) <-

OpenStudy (anonymous):

yes it is 2/7

OpenStudy (zzr0ck3r):

ok so now you have y=mx+b becomes y = (2/7)x+b use one of the points to solve for b 1 = (2/7)*4+b 1=(8/7)+b b = 1-(8/7) b = -(1/7) so y=(2/7)x-(1/7)

OpenStudy (jdoe0001):

after getting the slope 2/7 use the point-slope form => \(\bf y-y_1=m(x-x_1)\) and plug your values in now keep in mind you have to points, for the point-slope form either point is good, because both points are in the line, so you can use either \(\bf y-y_1=m(x-x_1)\) m = slope \(\bf (x_1, y_1)\) = either (4,1) or (-3,-1) then solve for "y", as shown above by zzr0ck3r